OpenStudy (anonymous):

10x-39+12x-9x+42-4x= 42 - 4x Solve each equation for x.

OpenStudy (anonymous):

First get all numbers on one side: For example, x + 6 = 2x -3 -6 -6 x = 2x -9 Second get all variables on one side: x = 2x -9 -2x -2x -x = -9 Third, divide each side by the coefficient of the variable (here, -1): -x = -9 -1 -1 x = 9

OpenStudy (anonymous):

this site will help you solve your question(s)

OpenStudy (anonymous):

Move all like terms to left hand side & all other to right hand side

OpenStudy (anonymous):

I did that and for some reason it came out wrong.. I know the answer is x=3 but I got something like 2.2.. Which isn't correct.

OpenStudy (anonymous):

can you tell me what is 10x + 12x - 9x - 4x = -------------?

OpenStudy (anonymous):

10x

OpenStudy (anonymous):

Please check it ..

OpenStudy (anonymous):

would it be 9x

OpenStudy (anonymous):

or -9x?

OpenStudy (anonymous):

Correct !! 9x so we will get 9x - 39 + 42 = 42 - 4x Right ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Now -39 +42 = ---------?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Correct !! so we have 9x + 3 = 42 - 4x Now move all like terms to left side & constant term to right side that is first add 4x to both sides so we get 9x + 3 + 4x = 42 - 4x +4x 9x+4x + 3 = 42 right ?

OpenStudy (anonymous):

Yes. so then it would be 12x + 3= 42

OpenStudy (anonymous):

Please check what is 9x + 4x = -------------?

OpenStudy (anonymous):

oops haha 13x

OpenStudy (anonymous):

so we have 13x + 3 = 42 To get rid of 3 from left side subtract 3 from both sides

OpenStudy (anonymous):

What do you get ?

OpenStudy (anonymous):

You get 13x=39 the divide right?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You are welcome :)